A goat is tied by a rope of length r metres to the corner of a square field with sides of length 100 m,. Find r if the goat is able to reach exactly half the grass in the field.


1
Expert's answer
2021-04-15T07:40:15-0400


Find the area of the field:


"S_1=100\cdot100=10000 m^2"

A goat is able to reach the area of a quarter of the circle:


"S_2 =\frac{1}{4}\pi r^2"

This area is equal to half the area of the field:


"\frac{1}{4}\pi r^2=5000""\pi r^2 =4\cdot 5000""\pi r^2 =20000""r^2 =\frac{20000}{\pi}""r=\sqrt{\frac{20000}{\pi}}=100\sqrt{\frac{2}{\pi}}\approx79.8 m"

Answer: The goat is able to reach exactly half the grass in the field if "r\approx79.8 m"